Express the given quantity in terms of the indicated variable.
The value (in cents) of the change in a purse that contains twice as many nickels as pennies, eight more dimes than nickels, and as many quarters as dimes and nickels combined; p = number of pennies
What I tried:
n=nickels, d=dimes, q=quarters
n=2p; d=2p+8; q=2p+(2p+8)
p+2p+(2p+8)+[2p+(2p+8)]
9p+16=wrong answer :(
: Express the given quantity in terms of the indicated variable.
The value (in cents) of the change in a purse that contains twice as many nickels as pennies, eight more dimes than nickels, and as many quarters as dimes and nickels combined; p = number of pennies
What I tried:
n=nickels, d=dimes, q=quarters
n=2p; d=2p+8; q=2p+(2p+8)
p+2p+(2p+8)+[2p+(2p+8)]
9p+16=wrong answer :(

Answer by checkley77(3624)

n=2p; d=2p+8; q=2p+(2p+8)
p+2p+(2p+8)+[2p+(2p+8)]
9p+16=wrong answer :(
YOU HAVE THE CORRECT ANSWER.
FOR EVERY PENNY YOU WILL HAVE:
1*9+16=25 OR:
1+2+10+12=25 COINS TOTAL.
FOR 2 PENNIES YOU WILL HAVE:
2*9+16=34
2+4+12+16=34 COINS TOTAL.
Question 156004: Express the given quantity in terms of the indicated variable.
The value (in cents) of the change in a purse that contains twice as many nickels as pennies, eight more dimes than nickels, and as many quarters as dimes and nickels combined; p = number of pennies
What I tried:
n=nickels, d=dimes, q=quarters
n=2p; d=2p+8; q=2p+(2p+8)
p+2p+(2p+8)+[2p+(2p+8)]
9p+16=wrong answer :(
: Express the given quantity in terms of the indicated variable.
The value (in cents) of the change in a purse that contains twice as many nickels as pennies, eight more dimes than nickels, and as many quarters as dimes and nickels combined; p = number of pennies
What I tried:
n=nickels, d=dimes, q=quarters
n=2p; d=2p+8; q=2p+(2p+8)
p+2p+(2p+8)+[2p+(2p+8)]
9p+16=wrong answer :(

Answer by scott8148(2758)

you found the number of coins __ the question asks for the value

try p+(2p)(5)+(2p+8)(10)+[2p+(2p+8)](25)